### v3

# Position IDs Matter: An Enhanced Position Layout for Efficient Context Compression in Large Language Models

Abstract
Using special tokens (e.g., gist, memory, or compressed tokens) to compress context information is a common practice for large language models (LLMs). However, existing approaches often neglect that position encodings inherently induce local inductive biases in models, causing the compression process to ignore holistic contextual dependencies. We propose Enhanced Position Layout (EPL), a simple yet effective method that improves the context compression capability of LLMs by only adjusting position IDs, the numerical identifiers that specify token positions. EPL minimizes the distance between context tokens and their corresponding special tokens and at the same time maintains the sequence order in position IDs between context tokens, special tokens, and the subsequent tokens. Integrating EPL into our best performing context compression model results in 1.9 ROUGE-1 F1 improvement on out-of-domain question answering datasets in average. When extended to multimodal scenarios, EPL brings an average accuracy gain of 2.6 to vision compression LLMs.  [^1] We will release the code upon acceptance.
Position IDs Matter: An Enhanced Position Layout for Efficient Context Compression in Large Language Models
Runsong Zhao1,*, Xin Liu1,*, Xinyu Liu1, Pengcheng Huang1, Chunyang Xiao, Tong Xiao1,2 and Jingbo Zhu1,2 1 NLP Lab, School of Computer Science and Engineering, Northeastern University, Shenyang, China 2 NiuTrans Research, Shenyang, China {fhe43694, chunyangx}@gmail.com {xiaotong, zhujingbo}@mail.neu.edu.com
11footnotetext: These authors contributed equally to this work.
1 Introduction
In Transformer Vaswani et al. (2017) architectures, special tokens have been widely adopted as compression carriers of contextual information across natural language processing Devlin et al. (2019); Liu et al. (2019); Bulatov et al. (2022); Ge et al. (2024); Li et al. (2024b) and computer vision Dosovitskiy et al. (2021); Ye et al. (2025). For context compression, so-called soft prompt methods Chang et al. (2024); Li et al. (2025) employ encoders to condense long contexts into few special tokens, enabling decoders to perform inference based on compressed representations rather than raw inputs, thereby significantly reducing memory consumption and inference latency in long-context scenarios Jiang et al. (2024); Xu et al. (2024). We illustrate typical soft prompt architectures in Figure 1 where special tokens are appended at the end of the context, intending to capture the context semantics via causal attention mechanism in LLMs.
The design ensures full context visibility for special tokens. However, we remark that in Transformer architectures position IDs do not need to coincide with physical token positions and the model’s perceived positional information is primarily determined by position IDs rather than physical token positions Vaswani et al. (2017). From this viewpoint, the local inductive biases introduced by position encodings Devlin et al. (2019); Vaswani et al. (2017); Su et al. (2023); Raffel et al. (2020); Press et al. (2022) weaken the efficacy of context compression under the default position layout (DPL), primarily because of the substantial distance between the special tokens and the context tokens, as illustrated in Figure 1 DPL. In this paper, we examine carefully position layout designs and propose Enhanced Position Layout (EPL) for soft prompt architectures, which comprises Uniform Position Layout (UPL) and Consistent Position Layout (CPL).
UPL redistributes special tokens’ position IDs to achieve uniform distribution in the context tokens’ position ID space, as exemplified in Figure 1. By uniformly assigning position IDs amongst context token position IDs, a priori, most context tokens would have corresponding special tokens close to them. We assume that such a prior helps the special tokens compress the context. We formalize such intuitions, demonstrating the optimality of the UPL in Section 3.2.1. During compression, because special tokens are inserted and text chunks are reorganized, the position IDs between context, special tokens and subsequent tokens (e.g. reconstructed tokens or subsequent tokens such as QA pairs) can become inconsistent compared to their original positions before compression. Our proposed CPL in Section 3.2.2 guarantees to maintain the position ID sequence order for different tokens, between different text chunks in their natural causal order.
We empirically apply EPL to two dominant context compression frameworks ICAE Ge et al. (2024) and 500xCompressor Li et al. (2024b). For the best model, on the autoencoding (AE) task, EPL yields a 1.8 BLEU gain and converges 9.7 times faster than DPL; on out-of-domain question answering (QA) tasks, EPL gives an average 1.9 ROUGE-1 F1 gain. When extending our application to multimodality with VoCo-LLaMA Ye et al. (2025), EPL yields an average 2.6 accuracy gain on multimodal benchmarks. The EPL improvement is consistent across base models of different scales. Further analysis shows that both UPL (which aims for better context compression) and CPL (which maintains causal sequence ordering) are essential for the final performance improvement across tasks. Finally, our UPL attention map visualization confirms the usefulness of our specified prior: UPL special tokens indeed focus more on tokens close to its assigned position IDs.
2 Background
2.1 Local Bias of Position Encodings
Transformer architectures Vaswani et al. (2017) compute contextual token embeddings through position-invariant self-attention. Since natural language semantics crucially depend on token order, various position encodings (PEs) have been proposed to inject positional awareness including Sinusoidal PE Vaswani et al. (2017), RoPE Su et al. (2023), Learnable PE Devlin et al. (2019), T5 Bias Raffel et al. (2020) and ALiBi Press et al. (2022), etc. All approaches share the inductive bias that adjacent tokens should correlate more strongly. Taking the PEs with the trigonometric encoding (e.g. Sinusoidal/RoPE) design as examples, the position embedding at a certain position is mostly similar to its neighbors and the similarity decay as the distance increase, see Appendix H.1 for more details. ALiBi enforce such inductive bias by applying distance-sensitive penalties to attention scores. We further show in Section 5.3 that Learnable PE and T5-bias learn similar local bias through pre-training on natural text.
Figure 1: Comparison of UPL and DPL. In prior work (DPL), memory tokens are assigned position IDs 7 and 8. Our method (UPL) allocates them to position IDs 2 and 5. Tokens with ID in close proximity tend to exhibit higher attention scores.
2.2 Position Layout
While the local inductive bias for PEs is well known, less is known about the position layout. For any given token sequence and model, we refer to position layout as the actual position ID sequence that is assigned by the model to the token sequence. Notably, in this work, we remark that the local inductive bias applies to the position layout, not the physical token positions.
This nuance is important but hardly noticeable because the default position layout (DPL) often coincides with the physical token positions as illustrated in DPL in Figure 1 (see Table 11 for more details on DPL). Such position layout is helpful for language modeling (LM) task since the task has been known to depend largely on its recent context Hu et al. (2024); Liu et al. (2024a). However, as LLMs are becoming a ubiquitous tool, we hypothesize that careful position layouts for some tasks can inject helpful inductive bias. In the rest of the paper, we focus on LLM compression tasks as our testbed.
3 Method
In Section 3.1, we review existing LLM compression frameworks and their DPLs; in Section 3.2 we describe our improved position layout.
3.1 Soft Prompt Methods
ICAE Ge et al. (2024) is a widely used encoder-decoder soft prompt method. Its encoder compresses long context into a few memory tokens, after which the decoder performs inference conditioned only on the memory tokens to achieve faster inference speed. Left of Figure 3 in Appendix illustrates this process through an example. ICAE can be trained through two stages: continued pretraining and fine-tuning. Continued pretraining trains on a combination of AutoEncoding (AE) tasks and Language Modeling (LM) tasks, which trains the LLM encoder so that the encoded memory tokens enable a frozen LLM to reconstruct losslessly the original context and at the same time predict the subsequent tokens following the context to maintain ICAE’s generation capability. Right of Figure 3 illustrates the AE training process. Fine-tuning further trains the ICAE encoder to adapt to real-word applications such as question answering (QA) that we use in this work. We review the AE and LM pretraining as well as QA finetuning in the next subsections. To handle arbitrarily long contexts, we adopt the multi-chunk version of ICAE, which divides any long context into chunks for independent compression and then aggregates the resulting memory tokens to represent the complete long context. 500xCompressor Ge et al. (2024) is similar to ICAE, with the difference of using the KV Cache of memory tokens as the compression carrier instead of the output of memory tokens.
3.1.1 Pretraining
During the pretraining stage, for a token sequence $X=\{x_{0},x_{1},\dots,x_{|X|-1}\}$, we take the first $p$ tokens as the context $X_{\text{context}}=\{x_{0},x_{1},\dots,x_{p-1}\}$ and the subsequent $|X|-p$ tokens as the completion $X_{\text{completion}}=\{x_{p},x_{p+1},\ldots,x_{|X|-1}\}$. The AE task only uses $X_{\text{context}}$, while the LM task leverages both $X_{\text{context}}$ and $X_{\text{completion}}$.
Compress
$X_{\text{context}}$ is partitioned into $k=\left\lceil p/L\right\rceil$ chunks (with chunk size $L$) where each chunk $S^{(i)}=\{x_{(i-1)L},x_{(i-1)L+1},\dots,x_{iL-1}\}$ is appended with a set of learnable memory tokens $M^{(i)}=\{m_{0}^{(i)},m_{1}^{(i)},\dots,m_{|M|-1}^{(i)}\}$. A LLM learns to encode each chunk into memory output tokens $\tilde{M}^{(i)}$ and key-value cache $KV^{(i)}$:
$\tilde{M}^{(i)},KV^{(i)}=\text{LLM}([S^{(i)};M^{(i)}]\mid\theta_{\text{LoRA}})$ (1)
where [;][;][ ; ] denotes concatenation along the sequence dimension. All $M^{(i)}$ share the same learnable parameters $M$, and $\theta_{\text{LoRA}}$ denotes a set of low-rank adapter Hu et al. (2022) parameters for the LLM. The final compressed representation is obtained by concatenating the results of each chunk: $\tilde{M}=[\tilde{M}^{(1)};\tilde{M}^{(2)};\dots;\tilde{M}^{(k)}]$ or $KV=[KV^{(1)};KV^{(2)};\dots;KV^{(k)}]$.
Pretraining
AE loss in ICAE is given by:
[…]
where [AE] is a learnable token prompting the frozen decoder to generate $X_{\text{context}}$ as reconstruction. Similar to the AE task, the loss for LM task in ICAE is given by:
[…]
where [LM] is a learnable token that prompts the LLM to perform completion. We employ a weighted loss function for joint training [^2] Unlike Ge et al. (2024)’s per-instance task allocation (AE with probability $\alpha$ /LM with $1-\alpha$), our joint training processes both tasks simultaneously through shared compressed representations, improving training efficiency by eliminating redundant context compression operations. :
$\mathcal{L}_{\text{pretrain}}=\alpha\mathcal{L}_{\text{AE}}+(1-\alpha)\mathcal{L}_{\text{LM}},\quad\alpha=0.5$
$k=2,L=510,|M|=102,r=5,|X|=2040,p=1020$
$S^{(1)}$ $S^{(2)}$
$x_{0}$ $x_{1}$ $x_{509}$ $m_{0}^{(1)}$ $m_{1}^{(1)}$ $m_{101}^{(1)}$ $x_{510}$ $x_{511}$ $x_{1019}$ $m_{0}^{(2)}$ $m_{1}^{(2)}$ $m_{101}^{(2)}$
DPL(ICAE/500x) 0 1 $509$ $510$ $511$ $611$ 0 1 $509$ $510$ $511$ $611$
EPL $1$ $2$ $510$ $3$ $8$ $508$ $511$ $512$ $1020$ $513$ $518$ $1018$
Table 1: Position layout of encoder for $S^{(1)}$ and $S^{(2)}$. More details see Appendix A, Table 12.
$k=2,L=510,|M|=102,r=5,|X|=2040,p=1020$
$\tilde{m}_{0}^{(1)}/KV_{0}^{(1)}$ $\tilde{m}_{101}^{(1)}/KV_{101}^{(1)}$ $\tilde{m}_{0}^{(2)}/KV_{0}^{(2)}$ $\tilde{m}_{101}^{(2)}/KV_{101}^{(2)}$ [AE] $x_{0}$ $x_{1}$ $x_{1019}$
DPL(ICAE) 00 $101$ $102$ $203$ $204$ $205$ $206$ $1224$
DPL(500x) $510$ $611$ $510$ $611$ $204$ $205$ $206$ $1224$
EPL $3$ $508$ $513$ $1018$ 00 $1$ $2$ $1020$
Table 2: Position layout example of decoder in AE Task. More details see Appendix A, Table 13.
$k=2,L=510,|M|=102,r=5,|X|=2040,p=1020$
$\tilde{m}_{0}^{(1)}/KV_{0}^{(1)}$ $\tilde{m}_{101}^{(1)}/KV_{101}^{(1)}$ $\tilde{m}_{0}^{(2)}/KV_{0}^{(2)}$ $\tilde{m}_{101}^{(2)}/KV_{101}^{(2)}$ [LM] $x_{1020}$ $x_{1021}$ $x_{2039}$
DPL(ICAE) 00 $101$ $102$ $203$ $204$ $205$ $206$ $1224$
DPL(500x) $510$ $611$ $510$ $611$ $204$ $205$ $206$ $1224$
EPL $3$ $508$ $513$ $1018$ $1020$ $1021$ $1022$ $2040$
Table 3: Position layout example of decoder in LM Task. More details see Appendix A, Table 14.
Position Layout
Recall that for encoding, the token sequence starts with text chunk $S^{(i)}$ followed by memory tokens $M^{(i)}$. For decoding, the token sequence starts with memory token outputs $\tilde{M}$ for ICAE or $KV$ for 500xCompressor followed by [LM] or [AE] and then subsequent tokens (i.e. $X_{\text{context}}$ or $X_{\text{completion}}$).
[…]
3.1.2 Fine-tuning
ICAE Ge et al. (2024) allows further fine-tuning to enhance downstream task performance by enabling memory tokens to learn to focus on context that most relate to the task; the training process is similar to the LM pretraining. Let each training instance consists of a triplet $(C,Q,A)$, where context $C$ is compressed into either $\tilde{M}$ (ICAE) or $KV$ (500xCompressor). The answer $A=\{a_{0},a_{1},...,a_{|A|-1}\}$ is generated conditioned on the compressed representation and the question $Q=\{q_{0},q_{1},...,q_{|Q|-1}\}$. The loss for QA task in ICAE is given by:
$\mathcal{L}_{\text{QA}}=-\log P\big{(}A\mid[\tilde{M};\texttt{[LM]};Q]\big{)}$ (4)
The QA loss for 500xCompressor is similar to Eq. (4), with the difference of conditioning on $KV$ instead of $\tilde{M}$.
Position Layout
The DPL for the QA task is similar to the DPL for the LM task and can be derived by replacing the $X_{\text{completion}}$ in the LM task with the concatenation $[Q;A]$. The resulting DPL for $[Q;A]$ coincides with their physical token positions. For the detailed DPL, see Table 15 in the appendix.
3.2 Enhanced Position Layout
For the DPL in soft prompt methods as described in Section 3.1, we identify two limitations:
Distant Memory Tokens
Memory tokens in soft prompt framework mainly aim to compress the context tokens so that the inference can be solely based on them to accelerate inference. However, the memory token DPL consists of a continuous range (i.e. $510$-$611$ in Table 1) and are all very distant from the context token’s range (i.e. 00-$509$). Given the local inductive bias of PEs that we briefly review in section 2.1, it would be advantageous to have memory token position IDs to be both close to context token position IDs and covering the context token ID range. In Section 3.2.1, we propose Uniform Position Layout (UPL) that has memory token position layout covering the context token ID range while achieving minimum ID distances between memory tokens and context tokens.
Inconsistent Layout
[…]
3.2.1 Uniform Position Layout
Algorithm 1 Generate Uniformly Distributed Compression Position IDs
1:Memory tokens count $|M|$, start/end IDs $v_{1}$, $v_{L}$
2:Uniform positions $U^{*}$
3:$r\leftarrow(v_{L}-v_{1}+1)/|M|$
4:$o\leftarrow\frac{r-1}{2}$
5:$U^{*}\leftarrow\text{torch.linspace}\bigl{(}v_{1}+o,v_{L}-o,\;|M|\bigr{)}$
[…]
Recall that for memory token position layout, we aim to achieve two objectives: (1) the memory token position ID should be close to context tokens (2) for any context token, there is some memory token(s) whose IDs are close to it. Figure 1 (b) Uniform Position Layout (UPL) illustrates such design, for any context token position ID, the nearest memory token position ID does not deviate more than 1, which sets contrast to DPL where the position ID of the first context token $x_{1}$ is far from all memory token position IDs. In the following, we formalize the desiderata to derive analytically the optimal position layout UPL.
Given a sequence of context token position IDs $V=\{v_{1},v_{2},\dots,v_{L}\}$, $v_{i}\in\mathbb{N}$ and $v_{i+1}-v_{i}=1$ $\forall i$, we aim to devise an algorithm to find position IDs for $|M|$ memory tokens $U=\{u_{1},u_{2},\dots,u_{|M|}\}$, $u_{j}\in\mathbb{N}$  [^5] Although non-integer position IDs are valid in RoPE Su et al. (2023), they have not been encountered during pre-training, making it difficult for the model to effectively utilize these non-integer position IDs., that minimize the following function:
$\max_{v_{i}\in V}\left(\min_{u_{j}\in U}|v_{i}-u_{j}|\right)$
where $\min_{u_{j}\in U}|v_{i}-u_{j}|$ represents the distance from the $i$-th context token to its nearest memory token, and $\max_{v_{i}\in V}$ takes the maximum of all minimum distances across context tokens.
The optimal solution divides $V$ evenly into $|M|$ groups, with each group containing at most $\left\lceil r\right\rceil$ tokens ($r=\frac{L}{|M|}$), and assigns each memory token the middle position of each group. In this case, the maximum distance from any context token to its nearest memory token is $\left\lfloor\frac{\left\lceil r\right\rceil}{2}\right\rfloor$ [^6] Note that any position layout will have its maximum distance $\geq\left\lfloor\frac{\left\lceil r\right\rceil}{2}\right\rfloor$, proving the optimality. . Intuitively, the solution spreads memory token position IDs uniformly in the range of context token position IDs to ensure that no context token position ID is too far away. We detail the memory token position layout algorithm in Algorithm 1 that we apply for each chunk to be compressed. Table 1 EPL row shows how the memory token position layout in UPL differs from DPL.
3.2.2 Consistent Position Layout
In this subsection, we propose consistent position layout (CPL) to ensure that the decoder position layout maintains the causal sequence order in position IDs between context tokens, [LM]/[AE], and the subsequent tokens. As shown in Table  2 and 3 EPL rows, we keep memory token position layout unchanged compared to its encoding stage.
[…]
4 Experimental Results
4.1 Experimental Setup
Data
For continued pretraining, we utilize the SlimPajama-6B Soboleva et al. (2023) corpus. To evaluate model fine-tuning performance, we use MRQA Fisch et al. (2019a) dataset as our testbed. The dataset contains evaluation on both in-domain scenarios where the validation dataset has its training counterpart used during training and out-of-domain scenarios. We report results from both settings but mainly discuss results for out-of-domain scenarios as it assesses more critically the soft prompt compression effectiveness.
Model Configuration
We evaluate our method on Llama-3.2-1B Grattafiori et al. (2024). For efficient adaptation, we apply LoRAHu et al. (2022) to the query and value projection matrices within the multi-head attention layers of the encoder. The LoRA rank is set to 128, and the LoRA alpha is set to 256. Following ICAE Ge et al. (2024), we do not train the decoder. In our default configuration, the number of memory tokens is $|M|=102$, the chunk size is $L=510$  [^8] As context often exceeds the chunk size, we extensively evaluate multi-chunk settings, contrary to ICAE and 500xCompressor., implying a $r=5$ compression ratio. All models are further pretrained and fine-tuned for 20k steps with a batch size of 16. Further hyperparameter details can be found in Appendix C, Table 8.
4.2 Fine-tuning Results
Following ICAE and 500xCompressor, we pretrain then fine-tune Llama-3.2-1B; we integrate our EPL changes described in Section 3.2 into different architectures respectively, then follow the same pretraining and finetuning steps. We evaluate the downstream performance using MRQA Fisch et al. (2019b) for different experiments. We assess the quality of the model’s answers using ROUGE-1 F1 Lin (2004) and Exact Match (EM) and report out-of-domain results in Table 4.
LM AE BioASQ DROP DuoRC RACE RE TQA Avg.
PPL BLEU F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
ICAE(DPL) 12.18 31.80 38.39 26.99 37.13 27.28 27.90 18.32 21.67 4.30 76.98 64.45 37.62 22.55 39.95 27.32
ICAE(EPL) 11.42 95.98 42.66 29.19 37.03 26.95 32.50 21.39 26.87 5.34 76.90 64.31 47.25 29.81 43.87 29.50
500x(DPL) 11.22 93.73 44.22 31.38 40.47 29.34 32.20 21.19 27.67 5.93 83.03 71.98 47.00 28.88 45.76 31.45
500x(EPL) 10.80 98.50 46.11 31.91 40.94 29.41 39.07 26.32 31.89 7.27 80.39 67.71 49.80 31.14 48.03 32.29
Table 4: Pretraining and fine-tuning results. For the pretraining results, we report perplexity for completion and BLEU-4 Papineni et al. (2002) score for reconstruction calculated on a held-out set of 1k examples. For the fine-tuning results, we report out-of-domain results, which includes 6 datasets: BioASQ Tsatsaronis et al. (2015), DROP Dua et al. (2019), DuoRC Saha et al. (2018), RACE Lai et al. (2017), Relation Extraction (RE) Levy et al. (2017), and TextbookQA (TQA) Kembhavi et al. (2017). F1 in all the tables is ROUGE-1 F1 Lin (2004).
For both ICAE and 500xCompressor, incorporating EPL significantly improves the performance. The average ROUGE-1 F1 improves from 39.95 to 43.87 for ICAE and improves from 45.76 to 48.03 for 500xCompressor. The improvement was observed for most domains, suggesting that the method is overall effective. We observe a similar improvement for in-domain settings (see Table 16 in the appendix).
4.3 Pretraining Results
Through pretraining LLMs have learned to compress context into memory tokens, allowing evaluation over memory tokens for its reconstruction and language modeling capability. The evaluation methodology is widely adopted for soft prompting Ge et al. (2024); Li et al. (2024a) and we expect EPL to bring a similar improvement to the fine-tuning settings since EPL incorporates useful prior to reconstruction and language modeling through its position layouts. The reconstruction quality and language modeling capability are evaluated using BLEU-4 Papineni et al. (2002) and perplexity (PPL), respectively [^9] Reconstruction texts are generated via greedy search..
Table 4 confirms the EPL improvement. For both ICAE and 500xCompressor architectures, we observe better language modeling capability with lower perplexity as well as better reconstruction capability with higher BLEU. The improvement is more significant with the weaker ICAE model but significant for both architectures.
4.4 Applications to Multimodal Models
As EPL can be applied to all applications that compress context into special tokens, in this subsection, we showcase its application in multimodaility. We follow VoCo-LLaMA Ye et al. (2025) for visual question-answering tasks. Given a triplet $(I,Q,A)$, the model encodes image $I$ into a sequence of 576 visual tokens $V^{t}=\{vt_{0},vt_{1},...,vt_{575}\}$, and subsequently compresses $V^{t}$ into the KV values of Vision Compression (VoCo) tokens. VoCo-LLaMA adopts a single training stage akin to fine-tuning stage in 500xCompressor and employs a single-forward via an attention mask (see Figure 8 for the attention mask detail) to prevent $Q$ and $A$ from directly accessing $V^{t}$. VoCo-LLaMA uses DPL and we follow its experimental setup to examine the effect of changing DPL to EPL. The position layout changes are illustrated on top of Figure 8.
We evaluate VoCo-LLaMA with 128 VoCo tokens (i.e. 4.5x compression ratio) and report performance on multimodal benchmarks. As shown in Table 5, VoCo-LLaMA combined with EPL significantly outperforms both its DPL counterpart from our reproduction and the results reported by Ye et al. (2025). We observe improvement across all three evaluated tasks, validating again the universality of EPL.
SQA(OOD) MMB(OOD) VQAv2(ID)
DPL* - 61.0 76.9
DPL 64.9 59.9 76.9
EPL 66.5 64.6 78.4
Table 5: Results of VoCo-LLaMA on SQA Lu et al. (2022), MMB Liu et al. (2024b), and VQAv2 Goyal et al. (2017). Results marked with * are from  Ye et al. (2025).
4.5 Ablation Studies
LM AE Out-of-Domain
PPL BLEU F1 EM
ICAE-1B 12.18 31.80 39.95 27.31
+UPL 11.50 72.35 43.67 29.11
+UPL & CPL (i.e. EPL) 11.42 95.98 43.87 29.50
500x-1B 11.22 93.73 45.76 31.45
+UPL 10.94 95.86 46.83 31.64
+UPL & CPL (i.e. EPL) 10.80 98.50 48.03 32.29
Table 6: Ablation study on EPL integration. +UPL applies only UPL to encoder; UPL & CPL (i.e. EPL) applies both UPL and CPL. Reported values are averages. More ablation results in Tables 18 and 19.
We conduct ablation studies to analyze the independent effects of UPL and CPL in EPL (see Section 3.2 for detailed description of UPL and CPL). LM perplexity and AE BLEU are measured under the same experimental settings as section 4.3 while out-of-domain performance is measured same as section 4.2. We show the results in Table 6 which confirm that:
UPL improves performance by injecting compression prior
For all testing cases (AE and LM during pretraining and MRQA out-of-domain performance for fine-tuning) over all architectures (ICAE and 500xCompressor), UPL improves performance compared to its counterparts. We note that the improvement is more significant for the weaker ICAE model. For example, the MRQA out-of-domain ROUGE-1 F1 improves 3.72 while the improvement was 1.07 for 500xCompressor. The results confirm that by carefully designing the memory position layout for compression tasks, the memory tokens obtain useful compression inductive prior, improves the compression efficiency (i.e. AE task performance) and consequently downstream task performance as shown by MRQA out-of-domain performance.
CPL improves performance by maintaining token sequential orders
[…]
4.6 Scalability Results
We conduct experiments at 3B and 8B scales from the same model family to verify method scalability. Table 7 shows that EPL achieves performance improvements at all scales and the improvement does not attenuate with larger scale. For example, at 8B scale, the MRQA ROUGE-1 F1 increased 14.03 for ICAE and increased 1.89 for 500xCompressor, which are similar to the improvement observed in 1B case in Table 4, holding promise for the method’s effectiveness on large-scale language models. More detailed results on MRQA can be found in Table 16 and 17.
LM AE Out-of-Domain
PPL BLEU F1 EM
Llama-3.2-3B
ICAE(DPL) 10.33 48.12 42.49 29.87
ICAE(EPL) 9.07 97.05 55.03 38.54
500x(DPL) 9.44 96.86 51.37 36.58
500x(EPL) 8.68 99.34 57.71 40.49
Llama-3.1-8B
ICAE(DPL) 9.22 49.08 43.09 30.01
ICAE(EPL) 7.61 98.82 57.12 39.76
500x(DPL) 7.61 97.68 57.42 40.32
500x(EPL) 7.40 99.49 59.31 41.45
Table 7: Results at Llama-3.2-3B Grattafiori et al. (2024), and Llama-3.1-8B. Reported values are averages.
5 Analysis
5.1 Training Curve
During pretraining, we observe that the adoption of EPL significantly accelerates the convergence speed of the AE loss. In the 8B-500xCompressor setting, for example, EPL reduces the training steps required to achieve an AE loss of 0.01 from 9.7k to 1.0k steps, while effectively mitigating AE loss fluctuations. This suggests that the prior information that EPL incorporates is well-suited for AE tasks. For more discussion, see Appendix E.
5.2 Attention Visualization
To verify whether the performance improvement we observe previously is indeed due to more suitable attention patterns for compression tasks idealized in Figure 1, we visualize the summed attention matrices of all attention heads in the second and final layers of the ICAE-3B model (Figure 6 and Figure 7) after the model has been fine-tuned on MRQA tasks.
For the second layer, we observe that the memory token under UPL attends to its surrounding [^10] In terms of position IDs, not physical token positions. context tokens, forming a slope (central top of Figure 6), contrast to DPL that only exhibits self-attention among memory tokens, showing that the attention adheres to our specified prior through EPL even after learning [^11] The angle of the slope is around $11^{\circ}$ ($\arctan(1/\textbf{5})=11.3^{\circ}$ at a compression rate of 5), which is consistent with our UPL conception illustrated in Figure 1. . In the final layer (Figure 7), UPL maintains the slope pattern, while DPL overcomes the position encoding resistance and attends to distant context tokens through learning without a clear pattern.
5.3 Learnable PEs
Our results rely on local bias of PEs that are explicitly defined in  Vaswani et al. (2017); Su et al. (2023); Press et al. (2022). We examine in this section empirically the properties of learnable position encodings Devlin et al. (2019); Raffel et al. (2020).
Figure 9 shows the cosine similarity of BERT’s Devlin et al. (2019) position embeddings where we observe significantly higher similarity between adjacent position embeddings compared to distant tokens. Similarly, The bias values of T5 Bias Raffel et al. (2020) decay with increasing relative distance as shown in Figure 2, indicating that attention is stronger amongst close tokens.
The BERT’s [CLS] token (position ID 0) shows a special pattern as shown in Figure 10: contrast to other tokens, its cosine similarities are not higher with its adjacent tokens. Given [CLS] token behaves like a compressed memory token trained using next sentence prediction, the phenomenon motivates our current work, suggesting that different priors should be given to special tokens for best performance.
6 Related Work
Soft Prompt Methods
GIST Mu et al. (2023) trained LLMs with modified attention mechanisms (similar to Figure 8) to compress prompt information into a few gist tokens. AutoCompressor Chevalier et al. (2023) trained an LLM to recursively compress long prompts by combining compressed tokens with new sub-prompts in each iteration, ultimately collecting all compressed tokens to form a compact representation. ICAE Ge et al. (2024) introduced an AE task enabling LLMs to pre-train compression capabilities on large-scale corpora, requiring only minimal parameter tuning for the encoder while freezing the decoder. 500xCompressor Li et al. (2024b) built upon ICAE by changing the information carrier from memory token outputs to memory tokens’ KV values. UniICL Gao et al. (2024b) and SelfCP Gao et al. (2024a) freeze both the encoder and decoder, training only a connector module to transform the encoder’s output memory tokens into decoder inputs. VoCo-LLaMA Ye et al. (2025) is the first to use LLMs for compressing visual tokens. Similar to GIST, it compresses visual token information into VoCo tokens through modified attention masks, outperforming methods like Q-Former Li et al. (2023) and average pooling with linear projection Li et al. (2024a). None of these methods discussed the impact of position layout and our EPL can be applied to all these soft prompt methods.
Position Layout
Although we are not aware of position layout work in the compression domain, we find related work in multimodality by considering memory tokens as another modality. Due to images being two-dimensional and text being one-dimensional, handling mixed image-text positional layout remains an open question. Current mainstream approaches in multimodality flatten 2D images into 1D sequences Liu et al. (2023b, a); Bavishi et al. (2023); Lu et al. (2024); Sun et al. (2024). Seeking a more elegant solution, Su (2024) proposed RoPE-Tie by placing visual and text tokens along the diagonal $(y=x)$ in 2D space. Although Su (2024) does not thoroughly validate their design, as the approach maintains the sequence order between modalities and the internal locality of images and text respectively, our empirical results suggest that the design can results in better performance. Qwen2-VL Wang et al. (2024) adopt similar designs in video domains.
7 Conclusion
We examine position layout, an understudied topic in context compression, and propose EPL for soft prompt methods. Our proposed position layout improves over default position layout by bringing memory tokens close to its context and at the same time maintains the logical sequence token ordering amongst context, memory tokens and the subsequent tokens. Extensive experiments demonstrate that EPL improves the compression efficiency and downstream task performance over different architectures across different modalities.
Given LLM’s ubiquitous usage, we believe that carefully examining position layout will give fruits to problems beyond context compression. We hope the success of EPL fosters research in this under explored area.
8 Limitations
Throughout our experiments, we have tested and confirmed EPL effectiveness under 5x compression ratio for text reconstruction and QA tasks and 4.5x compression ratio for visual QA. There remains question if our method is still effective at high compression ratio. We don’t have a positive outlook on this question.
First, EPL, requiring memory tokens to achieve uniform coverage across the context, intrinsically aligns with lossless compression scenarios (i.e., autoencoding tasks) which assumes that “all tokens in the context are equally important”; however, under high compression ratio (lossy compression), with information carrier capacity being limited, focus should be placed on important tokens, which violates our “equally important” assumption. Empirically, Figure 10 shows that the BERT’s [CLS] would have higher attention to some tokens without a clear pattern. Our preliminary results in Table 10 on VoCo-LLaMA also shows that EPL does not demonstrate any significant gains when used under high compression ratio (288x).
The analysis suggests that high ratio lossy compression may require compression mechanism that goes beyond our current work. At higher level, this suggests when adapting position layout methods to different application scenarios, the success can highly depend on whether the conceived layout captures the underlying prior characteristics of specific tasks.
Appendix A Detailed Position Layout
Detailed position layouts are provided in this section. Table 11 illustrates the position layout for LLMs without compression under standard conditions, reflecting natural language priors. Table 12 details the encoder’s position layout. For the decoder, position layouts are described in Tables 13, 14, and 15, with slight variations depending on the specific task (AE, LM, or QA). Each of these tables (Tables 12, 13, 14, and 15) includes a formula at the top representing the generalized layout. Below the formula, an illustrative example is given using a context length $|C|=p=1020$, chunk size $L=510$, compression ratio $r=5$, total sequence length $|X|=2040$, question length $|Q|=50$, and answer length $|A|=5$.
Appendix B Overview of ICAE
ICAE Ge et al. (2024) is an autoencoder framework to compress long contexts into short compact memory slots. The method operates by concatenating designated memory tokens to the end of the input sequence before an encoder processes the entire combined sequence. Subsequently, a decoder reconstructs the original sequence using only the information contained within the memory tokens. ICAE is trained in two main phases. It is first pretrained on massive text data using a combination of autoencoding and language modeling objectives, enabling it to generate memory slots that represent the original context. Following pretraining, the model is fine-tuned on instruction data for the purpose of producing desirable responses to various prompts. An overview of the ICAE framework is shown in Figure 3.
Appendix C Hyperparameters
For the 1B and 3B models, we perform continued pre-training on sequences with lengths $|X|$ ranging from 510 to 2040. For the 8B model, the input sequence length $|X|$ ranges from 510 to 4080 during continued pre-training. We take the first $p=\left\lfloor|X|/2\right\rfloor$ tokens as the context. Additional hyperparameters are listed in Table 8.
Hyperparameter Value
Optimizer AdamW
Betas (0.9, 0.95)
Weight decay 0.1
Learning rate 1e-4 (pretrain)
5e-5 (fine-tuning)
Scheduler Constant
Batch size 16
Warmup 300
Training steps 20k (pretrain)
20k (fine-tuning)
Clip norm 2.0
Table 8: Hyperparameters for training
Appendix D Detailed Results
In this section, we provide detailed results. Considering the potential risk of data leakage where LLMs may have encountered context information from evaluation datasets during the pretraining phaseLi et al. (2024b), we also report on NoContext and FullContext settings where NoContext performs inference solely based on the question and FullContext utilize the complete context and the question for inference. In both cases, we only train [AE] and [LM] tokens to guide the model executing corresponding tasks.
Table 9 fully presents the performance of ICAE and 500xCompressor of different scales on LM tasks and AE tasks. Table 16 and Table 17 respectively show all results of these models in-domain and out-of-domain in MRQA. Ablation results in-domain and out-of-domain in MRQA are presented in Table 18 and Table 19, respectively. Table 10 shows all results of VoCo-LLaMA on multimodal benchmarks.
PPL(AE) PPL(LM) BLEU(AE)
Llama-3.2-1B
NoContext 11.56 13.25 0.00
ICAE(DPL) 1.40 12.18 31.80
ICAE(EPL) 1.04 11.42 95.98
500x(DPL) 1.04 11.22 93.73
500x(EPL) 1.01 10.80 98.50
FullContext 1.02 9.90 33.49
Llama-3.2-3B
NoContext 9.58 11.02 0.00
ICAE(DPL) 1.49 10.33 48.12
ICAE(EPL) 1.02 9.07 97.05
500x(DPL) 1.02 9.44 96.86
500x(EPL) 1.00 8.68 99.34
FullContext 1.06 8.25 60.16
Llama-3.1-8B
NoContext 7.79 9.02 0.00
ICAE(DPL) 1.58 9.22 49.08
ICAE(EPL) 1.00 7.61 98.82
500x(DPL) 1.01 7.61 97.68
500x(EPL) 1.00 7.40 99.49
FullContext 1.00 7.30 98.00
Table 9: Perplexity of $X_{\text{context}}$ and $X_{\text{completion}}$, and BLEU score for the reconstruction quality of $X_{\text{context}}$, calculated on a held-out set of 1k examples.
voco_num SQA(OOD) MMB(OOD) VQAv2(ID)
Lower Bound* 1 60.7 22.3 41.2
DPL* 2 - 60.1 73.5
DPL 2 66.1 61.5 73.7
EPL 2 67.6 61.4 69.5
DPL* 128 - 61.0 76.9
DPL 128 64.9 59.9 76.9
EPL 128 66.5 64.6 78.4
Upper Bound* 576 66.5 64.0 77.7
Table 10: Results of VoCo-LLaMA on SQA Lu et al. (2022), MMB Liu et al. (2024b), and VQAv2 Goyal et al. (2017) from our experiments. Results marked with * are from  Ye et al. (2025).
Appendix E Training Curves
This section provides training curves. Figure 4 and Figure 5 respectively show the training curves of ICAE and 500xCompressor of different scales during pretraining. For AE loss, it can be observed that ICAE(DPL) struggles to decrease to 0, while 500xCompressor(DPL) requires a period of oscillation before converging near 0. When UPL is applied, their AE loss rapidly converges to around 0.
Appendix F VoCo-LLaMA
VoCo-LLaMA Ye et al. (2025) employs a modified attention mask which restricts text tokens from attending to vision tokens. Figure 8 illustrates this mask, and we have indicated its position layout at the top of the figure.
Appendix G Attention Visualization
Attention maps are presented in this section. Due to the low attention values of memory tokens in the first layer (which appear almost empty in the figure), we present the attention map of the second layer. See Figure 6. We provide a magnified view of the self-attention of memory tokens and their attention to other tokens. It can be observed that in the second layer, memory tokens in DPL only attend to themselves, while in UPL, they are able to attend to the entire context. Additionally, we also present the attention map of the last layer, shown in Figure 7. We use grey dashed lines to indicate the special attention pattern of UPL.
Appendix H Local bias of Position Encodings
H.1 Sinusoidal Position Encoding
The sinusoidal position encoding Vaswani et al. (2017); Su et al. (2023) is given by:
$PE_{(pos,2i)}=\sin\left(\frac{pos}{10000^{2i/d_{model}}}\right)$
$PE_{(pos,2i+1)}=\cos\left(\frac{pos}{10000^{2i/d_{model}}}\right)$
Consider two nearby positions, $pos$ and $pos+\delta$, where $\delta$ is a small number.
For any dimension $i$, the argument to the sine/cosine function changes from $\frac{pos}{10000^{2i/d_{model}}}$ to $\frac{pos+\delta}{10000^{2i/d_{model}}}$. The change in the argument is $\frac{\delta}{10000^{2i/d_{model}}}$.
For small $\delta$, this change in the argument is small for all dimensions.
Since sine and cosine functions are continuous, a small change in their input argument results in a small change in their output value.
$\sin(x+\epsilon)\approx\sin(x)\quad\text{for small }\epsilon$
$\cos(x+\epsilon)\approx\cos(x)\quad\text{for small }\epsilon$
Therefore, each dimension of $PE_{(pos+\delta)}$ is very close to the corresponding dimension of $PE_{(pos)}$.
This means the entire vector $PE_{(pos+\delta)}$ is very similar to $PE_{(pos)}$ when $\delta$ is small.
This property injects a local inductive bias.
H.2 Learnable Position Encodings
In this section, we present figures related to learnable position embeddings. Figure 9 illustrates the cosine similarity between different positions of BERT Devlin et al. (2019). It can be observed that the cosine similarity between nearby positions is significantly high. To illustrate the special behavior of the [CLS] token’s position embedding, we show in Figure 10 the cosine similarity of position 0 (i.e., the position ID of [CLS]), position 100, position 200, and position 300 with other positions. Figure 2 illustrates that as the relative distance increases, the learnable bias added by T5 Raffel et al. (2020) to the attention scores decreases.
Figure 2: T5 bias.
$x_{0}$ $x_{1}$ $x_{|X|-1}$ $y_{0}$ $y_{1}$ $y_{|Y|-1}$
Default Position ID 00 $1$ $|X|-1$ $|X|$ $|X|+1$ $|X|+|Y|-1$
$KV(x_{0})$ $KV(x_{1})$ $KV(x_{|X|-1})$ $y_{0}$ $y_{1}$ $y_{|Y|-1}$
Default Position ID 0 1 $|X|-1$ $|X|$ $|X|+1$ $|X|+|Y|-1$
Table 11: Default position layout of Transformers.
$x_{(i-1)L}$ $x_{(i-1)L+1}$ $x_{iL-1}$ $m_{0}$ $m_{1}$ $m_{|M|-1}$
DPL 0 1 $L-1$ $L$ $L+1$ $L+|M|-1$
EPL $(i-1)L+1$ $(i-1)L+2$ $iL$ $\left\lfloor b\right\rceil$ $\left\lfloor b+r\right\rceil$ $\left\lfloor(b+(|M|-1)r\right\rceil$
$i=1,L=510,|M|=102,r=5$
$x_{0}$ $x_{1}$ $x_{509}$ $m_{0}$ $m_{1}$ $m_{101}$
DPL 0 1 $509$ $510$ $511$ $611$
EPL $1$ $2$ $510$ $3$ $8$ $508$
$i=2,L=510,|M|=102,r=5$
$x_{510}$ $x_{511}$ $x_{1019}$ $m_{0}$ $m_{1}$ $m_{101}$
DPL 0 1 $509$ $510$ $511$ $611$
EPL $511$ $512$ $1020$ $513$ $518$ $1018$
[…]
$\tilde{m}_{0}^{(1)}/KV_{0}^{(1)}$ $\tilde{m}_{1}^{(1)}/KV_{1}^{(1)}$ $\tilde{m}_{|M|-1}^{(k)}/KV_{|M|-1}^{(k)}$ [AE] $x_{0}$ $x_{1}$ $x_{p-1}$
DPL(ICAE) 0 1 $k|M|-1$ $k|M|$ $k|M|+1$ $k|M|+2$ $k|M|+p$
DPL(500x) $L$ $L+1$ $L+|M|-1$ $k|M|$ $k|M|+1$ $k|M|+2$ $k|M|+p$
EPL $\left\lfloor b\right\rceil$ $\left\lfloor b+r\right\rceil$ $\left\lfloor b+(k|M|-1)r\right\rceil$ 00 $1$ $2$ $p$
$k=2,L=510,|M|=102,r=5,p=1020$
$\tilde{m}_{0}^{(1)}/KV_{0}^{(1)}$ $\tilde{m}_{1}^{(1)}/KV_{1}^{(1)}$ $\tilde{m}_{101}^{(2)}/KV_{101}^{(2)}$ [AE] $x_{0}$ $x_{1}$ $x_{1019}$
DPL(ICAE) 00 $1$ $203$ $204$ $205$ $206$ $1224$
DPL(500x) $510$ $511$ $611$ $204$ $205$ $206$ $1224$
EPL $3$ $8$ $1018$ 00 $1$ $2$ $1020$
[…]
$\tilde{m}_{0}^{(1)}/KV_{0}^{(1)}$ $\tilde{m}_{1}^{(1)}/KV_{1}^{(1)}$ $\tilde{m}_{|M|-1}^{(k)}/KV_{|M|-1}^{(k)}$ [LM] $x_{p}$ $x_{p+1}$ $x_{|X|-1}$
DPL(ICAE) 0 1 $k|M|-1$ $k|M|$ $k|M|+1$ $k|M|+2$ $k|M|+|X|-p$
DPL(500x) $L$ $L+1$ $L+|M|-1$ $k|M|$ $k|M|+1$ $k|M|+2$ $k|M|+|X|-p$
EPL $\left\lfloor b\right\rceil$ $\left\lfloor b+r\right\rceil$ $\left\lfloor b+(k|M|-1)r\right\rceil$ $p$ $p+1$ $p+2$ $|X|$
$k=2,L=510,|M|=102,r=5,|X|=2040,p=1020$
$\tilde{m}_{0}^{(1)}/KV_{0}^{(1)}$ $\tilde{m}_{1}^{(1)}/KV_{1}^{(1)}$ $\tilde{m}_{101}^{(2)}/KV_{101}^{(2)}$ [LM] $x_{1020}$ $x_{1021}$ $x_{2039}$
DPL(ICAE) 00 $1$ $203$ $204$ $205$ $206$ $1224$
DPL(500x) $510$ $511$ $611$ $204$ $205$ $206$ $1224$
EPL $3$ $8$ $1018$ $1020$ $1021$ $1022$ $2040$
[…]
$\tilde{m}_{0}^{(1)}/KV_{0}^{(1)}$ $\tilde{m}_{1}^{(1)}/KV_{1}^{(1)}$ $\tilde{m}_{|M|-1}^{(k)}/KV_{|M|-1}^{(k)}$ [LM] $q_{0}$ $q_{|Q|-1}$ $a_{0}$ $a_{|A|-1}$
DPL(ICAE) 0 1 $k|M|-1$ $k|M|$ $k|M|+1$ $t$ $t+1$ $t+|A|$
DPL(500x) $L$ $L+1$ $L+|M|-1$ $k|M|$ $k|M|+1$ $t$ $t+1$ $t+|A|$
EPL $\left\lfloor b\right\rceil$ $\left\lfloor b+r\right\rceil$ $\left\lfloor b+(k|M|-1)r\right\rceil$ $|C|$ $|C|+1$ $t^{\prime}$ $t^{\prime}+1$ $t^{\prime}+|A|$
$k=2,L=510,|M|=102,r=5,|C|=1020,|Q|=50,|A|=5$
$\tilde{m}_{0}^{(1)}/KV_{0}^{(1)}$ $\tilde{m}_{1}^{(1)}/KV_{1}^{(1)}$ $\tilde{m}_{101}^{(2)}/KV_{101}^{(2)}$ [LM] $q_{0}$ $q_{49}$ $a_{0}$ $a_{4}$
DPL(ICAE) 00 $1$ $203$ $204$ $205$ $254$ $255$ $259$
DPL(500x) $510$ $511$ $611$ $204$ $205$ $254$ $255$ $259$
EPL $3$ $8$ $1018$ $1020$ $1021$ $1070$ $1071$ $1075$
[…]
Figure 3: Overview of the ICAE framework proposed by Ge et al. (2024). Source: ICAE official repository (CC0-1.0).
SQuAD NewsQA TriQA SearchQA HQA NQ Avg.
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
Llama-3.2-1B
NoContext 9.34 1.92 4.80 0.59 3.04 0.92 14.99 8.71 9.47 3.27 10.88 3.91 8.75 3.22
ICAE(DPL) 54.58 36.99 37.57 21.25 57.96 48.85 68.90 56.68 59.61 43.01 59.40 42.01 56.34 41.47
ICAE(EPL) 59.94 39.39 43.61 24.95 61.50 51.98 69.31 57.11 64.22 46.84 61.21 42.11 59.97 43.73
500x(DPL) 68.26 49.15 43.17 25.31 60.30 51.03 70.33 57.96 66.44 49.45 64.69 46.78 62.20 46.61
500x(EPL) 67.87 47.54 49.23 29.58 64.53 55.16 72.51 60.39 68.51 51.06 65.55 46.97 64.70 48.45
FullContext 58.72 39.20 38.72 16.62 31.82 24.42 49.15 36.27 53.22 39.08 53.47 36.12 47.51 31.95
Llama-3.2-3B
NoContext 15.65 7.09 6.15 1.26 8.68 6.38 42.54 31.83 16.31 9.02 16.43 8.18 17.63 10.63
ICAE(DPL) 58.95 41.53 39.70 23.12 60.49 51.64 66.35 53.81 58.68 42.55 58.55 41.94 57.12 42.43
ICAE(EPL) 73.82 53.67 58.02 36.51 71.22 62.26 71.41 59.90 72.61 55.77 70.19 51.33 69.55 53.24
500x(DPL) 68.28 50.15 50.98 32.15 68.10 59.54 74.16 62.11 70.63 53.57 67.58 50.19 66.62 51.28
500x(EPL) 77.74 58.38 61.19 41.62 71.62 62.66 74.32 62.67 75.06 58.52 72.03 54.07 71.99 56.32
FullContext 74.07 55.43 48.99 24.62 63.92 54.05 67.69 52.31 63.73 48.33 64.32 46.54 63.79 46.88
Llama-3.1-8B
NoContext 21.84 11.74 9.52 3.28 31.58 26.37 59.66 45.49 20.34 12.71 29.62 17.94 28.76 19.59
ICAE(DPL) 56.56 38.87 36.99 20.06 64.54 55.58 71.35 58.92 57.04 41.20 58.04 41.19 57.42 42.64
ICAE(EPL) 78.44 58.90 61.69 40.17 73.92 65.05 80.05 67.65 74.93 58.33 72.43 54.14 73.58 57.37
500x(DPL) 80.56 62.11 60.31 40.65 74.00 65.39 79.60 67.51 76.18 59.62 74.17 56.48 74.14 58.63
500x(EPL) 80.60 61.37 64.43 44.35 74.75 66.04 79.39 67.74 77.17 60.53 74.71 56.16 75.18 59.36
FullContext 80.53 61.97 60.24 40.05 72.65 63.28 76.54 61.99 73.07 57.19 72.25 54.01 72.55 56.42
Table 16: Results on the in-domain validation set, including six QA datasets: SQuAD Rajpurkar et al. (2016), NewsQA Trischler et al. (2017), TriviaQA (TriQA) Joshi et al. (2017), SearchQA Dunn et al. (2017), HotpotQA (HQA) Yang et al. (2018), and NaturalQuestions (NQ) Kwiatkowski et al. (2019).
BioASQ DROP DuoRC RACE RE TQA Avg.
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
Llama-3.2-1B
NoContext 10.63 4.52 17.72 10.18 3.93 0.53 6.58 0.30 11.74 3.70 20.50 9.78 11.85 4.83
ICAE(DPL) 38.39 26.99 37.13 27.28 27.90 18.32 21.67 4.30 76.98 64.45 37.62 22.55 39.95 27.32
ICAE(EPL) 42.66 29.19 37.03 26.95 32.50 21.39 26.87 5.34 76.90 64.31 47.25 29.81 43.87 29.50
500x(DPL) 44.22 31.38 40.47 29.34 32.20 21.19 27.67 5.93 83.03 71.98 47.00 28.88 45.76 31.45
500x(EPL) 46.11 31.91 40.94 29.41 39.07 26.32 31.89 7.27 80.39 67.71 49.80 31.14 48.03 32.29
FullContext 45.72 31.45 40.74 30.27 39.20 27.85 29.57 7.72 74.85 64.01 53.31 34.60 47.23 32.65
Llama-3.2-3B
NoContext 24.84 17.22 21.17 13.37 5.87 1.93 8.87 1.19 20.60 13.30 37.34 22.82 19.78 11.64
ICAE(DPL) 41.75 31.65 38.77 29.81 28.40 18.99 21.58 3.56 76.56 64.82 47.85 30.41 42.49 29.87
ICAE(EPL) 50.83 36.70 54.10 43.38 45.66 33.24 38.78 9.35 82.71 71.64 58.13 36.93 55.03 38.54
500x(DPL) 50.20 37.63 48.64 38.32 38.77 26.85 31.86 8.16 83.67 73.71 55.08 34.80 51.37 36.58
500x(EPL) 53.15 39.03 57.00 45.84 50.35 36.51 42.45 10.83 85.20 75.20 58.09 35.53 57.71 40.49
FullContext 59.20 42.82 50.30 35.53 39.09 27.51 36.72 9.20 81.73 73.27 67.50 43.65 55.76 38.66
Llama-3.2-8B
NoContext 43.16 33.64 25.60 18.43 6.15 2.33 7.98 1.63 33.89 25.27 50.27 32.67 27.84 19.00
ICAE(DPL) 47.42 33.98 35.98 27.41 22.54 13.99 21.48 4.60 77.58 65.98 53.55 34.13 43.09 30.01
ICAE(EPL) 53.21 37.90 58.83 47.50 47.26 33.44 40.65 9.64 84.35 73.91 58.39 36.13 57.12 39.76
500x(DPL) 51.50 37.70 59.85 47.64 47.95 34.44 41.17 10.39 85.86 75.85 58.22 35.93 57.42 40.32
500x(EPL) 54.80 39.43 62.46 50.70 51.33 37.38 43.40 11.13 85.77 75.10 58.11 35.00 59.31 41.45
FullContext 59.80 42.95 60.64 47.64 26.91 16.86 42.80 10.53 85.11 74.08 71.68 47.57 57.83 39.94
Table 17: Results on the out-of-domain validation set, including six QA datasets: BioASQ Tsatsaronis et al. (2015), DROP Dua et al. (2019), DuoRC Saha et al. (2018), RACE Lai et al. (2017), Relation Extraction (RE) Levy et al. (2017), and TextbookQA (TQA) Kembhavi et al. (2017).
SQuAD NewsQA TriQA SearchQA HQA NQ Avg.
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
Llama-3.2-1B
ICAE(DPL) 54.58 36.99 37.57 21.25 57.96 48.85 68.90 56.68 59.61 43.01 59.40 42.01 56.34 41.47
ICAE(UPL) 62.22 42.25 44.49 26.16 61.35 51.88 70.33 57.73 63.85 46.89 61.71 43.00 60.66 44.65
ICAE(EPL) 59.94 39.39 43.61 24.95 61.50 51.98 69.31 57.11 64.22 46.84 61.21 42.11 59.97 43.73
500x(DPL) 68.26 49.15 43.17 25.31 60.30 51.03 70.33 57.96 66.44 49.45 64.69 46.78 62.20 46.61
500x(UPL) 65.92 46.35 49.33 29.89 63.97 54.98 71.85 59.77 67.04 50.53 64.20 45.66 63.72 47.86
500x(EPL) 67.87 47.54 49.23 29.58 64.53 55.16 72.51 60.39 68.51 51.06 65.55 46.97 64.70 48.45
Table 18: Ablation results on the in-domain validation set, including six QA datasets: SQuAD Rajpurkar et al. (2016), NewsQA Trischler et al. (2017), TriviaQA (TriQA) Joshi et al. (2017), SearchQA Dunn et al. (2017), HotpotQA (HQA) Yang et al. (2018), and NaturalQuestions (NQ) Kwiatkowski et al. (2019).
BioASQ DROP DuoRC RACE RE TQA Avg.
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
Llama-3.2-1B
ICAE(DPL) 38.39 26.99 37.13 27.28 27.90 18.32 21.67 4.30 76.98 64.45 37.62 22.55 39.95 27.32
ICAE(UPL) 41.82 29.32 37.77 27.41 33.73 22.45 26.90 4.90 73.61 60.62 48.19 29.94 43.67 29.11
ICAE(EPL) 42.66 29.19 37.03 26.95 32.50 21.39 26.87 5.34 76.90 64.31 47.25 29.81 43.87 29.50
500x(DPL) 44.22 31.38 40.47 29.34 32.20 21.19 27.67 5.93 83.03 71.98 47.00 28.88 45.76 31.45
500x(UPL) 45.71 32.38 40.19 28.81 37.66 26.12 29.02 6.08 78.22 64.55 50.19 31.87 46.83 31.64
500x(EPL) 46.11 31.91 40.94 29.41 39.07 26.32 31.89 7.27 80.39 67.71 49.80 31.14 48.03 32.29
Table 19: Ablation results on the out-of-domain validation set, including six QA datasets: BioASQ Tsatsaronis et al. (2015), DROP Dua et al. (2019), DuoRC Saha et al. (2018), RACE Lai et al. (2017), Relation Extraction (RE) Levy et al. (2017), and TextbookQA (TQA) Kembhavi et al. (2017).
Figure 4: Training Loss of ICAE.
Figure 5: Training Loss of 500xCompressor.
Figure 6: Attention Matrix in the 2nd layer of ICAE after MRQA finetuning.
Figure 7: Attention Matrix in the last layer of ICAE after MRQA fine-tuning.
Figure 8: Attention mask and position layout of VoCo-LLaMA Ye et al. (2025). The Position IDs modified by EPL are marked in green (UPL) and blue (CPL) on top of the figure.
Figure 9: The cosine similarity of BERT’s positional encodings. To improve readability, we have removed results beyond the top 10 and zoomed in on the first 50 positions.
Figure 10: The cosine similarity between the positional encoding of the [CLS] token and other positions.